OpenStudy (anonymous):

If a circuit has four resistors connected in parallel with resistances of 20 Ohms each, what is the total resistance of the circuit? A. 20 ohms B. 5 ohms C. 80 ohms

OpenStudy (perl):

$$ \Large { \frac{1}{R_{Total}}= \frac{1}{R_{1}} +\frac{1}{R_{2}}+\frac{1}{R_{3}}+\frac{1}{R_{4}} } $$

OpenStudy (perl):

now you can plug in, but dont forget to find the reciprocal of the sum

OpenStudy (anonymous):

b

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

no the answer is 5...20/4= 5. Right @perl

OpenStudy (perl):

correct

OpenStudy (anonymous):

your awesome

OpenStudy (perl):

:)

OpenStudy (anonymous):

As more resistors are connected in a parallel circuit, the overall current in the circuit: A. increases. B. decreases. C. stays the same

OpenStudy (irishboy123):

adding more parallel resistors is completely analogous to increasing the cross sectional area of a single resistor. in each case, you increase the number of valence electrons that are available to be moved into the conduction band by the electric field and then shuffled around the circuit as current. the flip side is that you drain more power from the source when you do this. this conclusion is also reflected in the related equations as these empirical equations are driven by observation. R = 1/ ∑1/r so adding in r's always decreases the overall R.